OpenStudy (anonymous):

Rational Exponents again (x^1/3 / y^-2/3)^9

OpenStudy (jhannybean):

\[\large \left(\frac{x^{1/3}}{y^{-2/3}}\right)^9 \]

OpenStudy (jhannybean):

Remember, \[x^{-1} \iff \frac{1}{x}~ , ~ \frac{1}{x^{-1}} \iff x^1\]

OpenStudy (jhannybean):

So which of those two identities would you use for \(y^{-2/3}\)?

OpenStudy (anonymous):

1/y^2/3?

OpenStudy (jhannybean):

Not quite. The second identity claims this: \[\frac{1}{x^{-1}} = x^{1}\]This meant you can take the reciprocal of the function in the denominator of the fraction and raise it up to the numerator.

OpenStudy (anonymous):

oh wait the other one so i can get rid of the denominator

OpenStudy (jhannybean):

Yes \(\checkmark\)

OpenStudy (jhannybean):

So what would your equation look like within the parenthesis if you used the second identity on \(y^{-2/3}\)?

OpenStudy (anonymous):

(x^1/3 y^2/3)9

OpenStudy (anonymous):

so then its just cross multiply again?

OpenStudy (jhannybean):

Awesome. You're a quick learner, I like that

OpenStudy (jhannybean):

Don't call it cross multiplication, call it distribution. You're now distributing the exponent of 9 to both x and y terms.

OpenStudy (jhannybean):

So what is \[\frac{1}{3} \cdot 9 = ~? \]\[\frac{2}{3} \cdot 9 =~? \]

OpenStudy (anonymous):

so x would be x^3 and y would be y^6

OpenStudy (jhannybean):

\(\checkmark\)

OpenStudy (anonymous):

awesome thanks

OpenStudy (jhannybean):

No problem